Q.#1 Using Excel format:

A company is studying two different production methods to produce its product. Both production methods have the same life and the same maintenance and repair record.

  • Production Method 1: costs $120,000 and uses 20 gallons per hours of operation.
  • Production Method 2: costs $175,000 and uses 18 gallons per hours of operation at the same level of production.

Both production methods have four-year lives before any major overhaul is required and the 10% of initial its cost as a salvage value. The fuel currently costs $2.40

The fuel consumption is expected to increase 1gallon/hr. every year because of degrading engine efficiency and fuel costs is expected to increase at the rate of 5% per year.

Required:

Assume 5,000 hours of operation per year and a MARR of 12%. Use the AE criterion.

  1. Summarize the Net Cash Flow of Production Method 1 in table format.
  2. Summarize the Net Cash Flow of Production Method 2 in table format
Answer #1

Using NPV & PMT function in EXCEL

Cash Flow associated with Method 1

Method 1
Year Investment Annual Fuel consumption Salvage Net Cash Flow
0 -120000.00 -120000.00
1 -252000.00 -252000.00
2 -277830.00 -277830.00
3 -305613.00 -305613.00
4 -335479.73 12000.00 -323479.73
Annual worth -3,25,807.40

Cash Flow associated with Method 2

Method 2
Year Investment Annual Fuel consumption Salvage Net Cash Flow
0 -175000.00 -175000.00
1 -226800.00 -226800.00
2 -251370.00 -251370.00
3 -277830.00 -277830.00
4 -306307.58 17500.00 -288807.58
Annual worth -3,15,797.38

As Annual cost of method 2 is lower, it should be selected

Showing formula in excel

Method 1
Year Investment Annual Fuel consumption Salvage Net Cash Flow
0 -120000 =B28+C28+D28
1 =-20*5000*2.4*1.05^A29 =B29+C29+D29
2 =-21*5000*2.4*1.05^A30 =B30+C30+D30
3 =-22*5000*2.4*1.05^A31 =B31+C31+D31
4 =-23*5000*2.4*1.05^A32 =0.1*120000 =B32+C32+D32
Annual worth =PMT(12%,4,-(NPV(12%,E29:E32)+E28))
Method 2
Year Investment Annual Fuel consumption Salvage Net Cash Flow
0 -175000 =B37+C37+D37
1 =-18*5000*2.4*1.05^A38 =B38+C38+D38
2 =-19*5000*2.4*1.05^A39 =B39+C39+D39
3 =-20*5000*2.4*1.05^A40 =B40+C40+D40
4 =-21*5000*2.4*1.05^A41 =0.1*175000 =B41+C41+D41
Annual worth =PMT(12%,4,-(NPV(12%,E38:E41)+E37))